OpenStudy (anonymous):

lim x approaches a (A^3-X^3)/(A-X)

OpenStudy (anonymous):

use a^3-b^3= (a-b) (a^2+ab+b^2)

OpenStudy (anonymous):

factor the numerator (difference of cubes)... simplify, then to find the limit just do a direct substitution.

OpenStudy (anonymous):

and simplify the numerator to cancel common terms with denominator

OpenStudy (anonymous):

I get 3a^2.Just wanted to check my answer

OpenStudy (anonymous):

can you see if you get the same result

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Thanks one more question Describe the influence of the three parameters, A, k, and n, on the shape of the graph of the Hill function y=(AX^n)/(x^n+k^n)